I measured the inside diameter of a cylinder as 3.42 cm and the outside diameter as 4.60 cm.
(i) calculate the maximum thickness and the minimum thickness of the cylinder wall
(ii) state the value of the thickness with its limits of error.
(iii) write down the thickness to an appropriate number of significant figures.
1
Expert's answer
2019-09-26T14:48:28-0400

(i) for maximum thick we will take upper certainity of outside diameter and lower certainity of inside diameter and vice-versa for minimum diameter

So, Maximum thickness = "(4.60+0.01)-(3.42-0.01)=1.20\ cm"

Minimum thickness =,"(4.60-0.01)-(3.42+0.01)=1.16\ cm"


(ii) value of thickness = "(4.60\pm0.01)-(3.42\pm0.01)=(1.18\pm0.02)\ cm"


(iii) thickness = "4.60-3.42=1.18\ cm"